OpenStudy (anonymous):

The area of the cube is increasing at a rate 4 cm^2/s. How fast is the volume of the cube increasing when the edge length is 10 cm?

OpenStudy (anonymous):

Given: \[\frac{ dA }{ dt }=4cm ^{2}/s\] \[V=L*W*H\] \[\frac{ dV }{ dt }=WH*\frac{ dL }{ dt }+LH*\frac{ dW }{ dt }+LW*\frac{ dH }{ dt }\] Since this is a cube all the sides are equal. \[\frac{ dV }{ dt }=100(\frac{ dL }{ dt }+\frac{ dW }{ dt }+\frac{ dH }{ dt })\] \[A=6a^2, \frac{ dA }{ dt }=12a, a=\frac{ 4 }{ 12 }\] \[\frac{ dV }{ dt }=100(\frac{ 4 }{ 12 })=\frac{ 100 }{ 3 }cm ^{3}/s\] Is this right?

OpenStudy (anonymous):

is it just me or r u guys also losing connection every few minutes

OpenStudy (freckles):

yes I am :(

OpenStudy (freckles):

ok anyways a seems to be always changing w.r.t to time so a=a(t) where a is an edge of a cube \[A(t)=6 [a(t)])^2\] \[V(t)=[a(t)]^3\] we are given A'(t)=4

jimthompson5910 (jim_thompson5910):

s = side length of cube A = surface area of cube A = 6s^2 dA/dt = 12s*ds/dt 4 = 12s*ds/dt ds/dt = (4)/(12s) ds/dt = 1/(3s) V = volume of cube V = s^3 dV/dt = 3s^2*(ds/dt) dV/dt = 3s^2*(1/(3s)) dV/dt = (3s^2)/(3s) dV/dt = s dV/dt = 10

OpenStudy (freckles):

\[A'(t)=12 a(t) \cdot a'(t)\] \[12 a(t) \cdot a'(t)=4\] Solve for a'(t)\] we will later use a(t) is 10 after we take care of all the differentiating

OpenStudy (freckles):

nevermind @jim_thompson5910 beat me to it

OpenStudy (anonymous):

lol no worries

OpenStudy (anonymous):

ah ok I see

OpenStudy (anonymous):

Thank you guys!